OpenStudy (anonymous):

The table below represents the distance of a car from its destination as a function of time: x y 0 900 1 850 2 800 3 750

OpenStudy (anonymous):

Part A: What is the y-intercept of the function, and what does this tell you about the car? The car is traveling 50 miles per hour because you can see that the distance from it's destination is decreasing by 50. Need to check my answer

OpenStudy (anonymous):

@ninjasandtigers

OpenStudy (anonymous):

The distance traveled is decreasing. So I thnk he is traveling backward, but still it is 50mps yes

OpenStudy (anonymous):

OK thanks, But I need help with the rest

OpenStudy (anonymous):

Part B: Calculate the average rate of change of the function represented by the table between x = 1 to x = 3, and tell what the average rate represents.

OpenStudy (anonymous):

hmm so the formula for average rate of change would be

OpenStudy (anonymous):

It's slope, but I keep forgetting the formula -_-

OpenStudy (anonymous):

so it will be f(b)-f(a)/b-a

OpenStudy (anonymous):

\[\frac{ y _{2}-y _{1} }{ x _{2}-x}\]

OpenStudy (anonymous):

can you plug in the numbers for me?

OpenStudy (anonymous):

oh and yeah that slope formula could work out too

OpenStudy (anonymous):

that should be x_2-x_1

OpenStudy (anonymous):

Yes! Thanks @ninjasandtigers lol i knew the formula from when you told me earlier but I forgetting it already lol

OpenStudy (anonymous):

haha your welcome :)

OpenStudy (anonymous):

what would you get if you calculated?

OpenStudy (anonymous):

so 0 = x1 1 = x2 900 = y1 850 = y2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

850 - 900 = -50

OpenStudy (anonymous):

\[\frac{ 850-900 }{ 1-0}\]

OpenStudy (anonymous):

1 - 0 = 1

OpenStudy (anonymous):

-50 / 1 = -50

OpenStudy (anonymous):

thats it!!!!

OpenStudy (anonymous):

yes so that is your slope :)

OpenStudy (anonymous):

now plug in any set of coordinates into y and x

OpenStudy (anonymous):

Wait, but it says x = 1 TO x = 3? I should have used (3,750) right?

OpenStudy (anonymous):

the slope from (1,850) to (3,750) would be -100

OpenStudy (anonymous):

Ah ok, hold on let me rewrite the answer

OpenStudy (anonymous):

1 = x1 3 = x2 850 = y1 750 = y2 750 - 850 = -100 3 - 1 = 2 -100/2 = -50

OpenStudy (anonymous):

Part C: What would be the domain of this function if the car traveled the same rate until it reached its destination?

OpenStudy (anonymous):

did you already find the y intercept?

OpenStudy (anonymous):

OH WAIT lol no i didnt!

OpenStudy (anonymous):

ok so lets use (0,900)

OpenStudy (anonymous):

y=mx+b 900=-50(0)+b 900=b

OpenStudy (anonymous):

Okie, now what?

OpenStudy (anonymous):

ok so, y=-50x+900

OpenStudy (anonymous):

you don't know where the destination is, so it will be all real numbers

OpenStudy (anonymous):

ok but what is the domain?

OpenStudy (anonymous):

that is the domain, all real numbers, because you don't know where the destination is so the car could stop anywhere

OpenStudy (anonymous):

Ohh.. Okay, thank you so much! :)

OpenStudy (anonymous):

welcome :D